OpenStudy (anonymous):

A cell phone plan has a monthly cost that is shown in the table below. What is the correct statement regarding the average rate of change during the 40-minute time of talk? Total minutes of talk time: 0 10 20 30 40 Monthly cost of cell phone: $19.95 $20.45 $20.95 $21.45 $21.95 The average rate of change is $0.50, meaning that for each minute of talk time, the monthly bill increases by $0.50 The avaerage rate of change is $0.50, meaning that for every ten minutes of talk time, the bill increases by $0.50

OpenStudy (anonymous):

The average rate of change is $0.05, meaning that for each minute of talk time, the monthly bil increaes by $0.05 The avaerage rate of change is $0.05, meaning that for every ten minutes of talk time, the bill increases by $0.05

OpenStudy (anonymous):

I think it's the second of fourth one, but I'm not sure which one?

OpenStudy (tkhunny):

Try First Differences 0 10 20 30 40 10-0 = 10 20 - 10 = 10 30 - 20 = 10 40 - 30 = 10 They are all 10 minutes apart. $19.95 $20.45 $20.95 $21.45 $21.95 20.45 - 19.95 = 0.50 20.95 - 20.45 = 0.50 21.45 - 20.95 = 0.50 21.95 - 21.45 = 0.50 The all change by 50¢. Just looking at it, for each 10 minutes of talk time, the bill goes up 50¢. Dividing both by 10 suggests, for each 1 minute of talk time, the bill goes up 5¢. Just some hints on one method that might be used to analyze the situation.

OpenStudy (anonymous):

y=1/20x + 19.95

OpenStudy (tkhunny):

Oh, and get a better User ID. How do you plan to succeed when you talk about yourself like that?

OpenStudy (anonymous):

I think slope-intercept form represents this equation fine.

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

the slope of 1/20 represents 5cents per minute

OpenStudy (anonymous):

So it would be B? I just came up with something off the top of my head for my ID lol was planning on changing it soon haha

OpenStudy (anonymous):

I dont have your choices, therefore, i cannot tell you whether the answer is b or not.

OpenStudy (anonymous):

B is The average rate of change is $0.50, meaning that for every ten minutes of talk time, the monthly bill increases by $0.50

OpenStudy (anonymous):

It seems like it can be either B or C /:

OpenStudy (anonymous):

the average rate of change is indeed 50 cents every ten minutes, equal to 5 cents every minute.